On January 1, 2021, CLOY Company granted a five-year term loan on P1,500,000 on GOBLIN Company. If there were no possibility of credit losses, the coupon rate that CLOY Company would charge the borrower is 10% per annum, because of the borrower’s credit rating, CLOY Company estimates that there is a possibility the borrower might default on the payments and the expected credit losses are estimated at P20,000 per year over the five-year term. Accordingly, CLOY Company charges the borrower 12% coupon rate to reflect the yield on the instrument to include a return to cover those credit losses expected when the loan is first recognized.

1. Compute for the lifetime expected credit loss.

2. Compute for the 12-month expected credit loss.

3. Prepare the journal entry on initial recognition of the loan.

4. Prepare the entry assuming there is no significant deterioration of credit risk for the year ended 2021,

5. Prepare the entry assuming there is significant deterioration of credit risk for the year ended 2021.



1
Expert's answer
2021-06-07T08:14:02-0400

Solution:

1.). The lifetime expected credit loss:

Using the Discounted Cash flows approach:

The lifetime expected credit loss = "\sum \frac{Expected \;credit\; losses}{(1 + effective\; interest \;rate)}"


Expected losses = 20,000 per year


Effective interest rate = "12\%"

The lifetime expected credit loss = "\sum \frac{20,000}{(1 + 0.12)^{1} } + \frac{20,000}{(1 + 0.12)^{2} } + \frac{20,000}{(1 + 0.12)^{3} } + \frac{20,000}{(1 + 0.12)^{4} } + \frac{20,000}{(1 + 0.12)^{5} }"


= 17,857 + 15,944 + 14,236 + 12,710 + 11,349 = 72,096

The lifetime expected credit loss = 72,096


2.). The 12-month expected credit loss:

The 12-month expected credit loss= "\frac{20,000}{(1 + 0.12)^{1} } = 17,857"


The 12-month expected credit loss = 17,857


3.). The journal entry for initial recognition of the loan:

On initial recognition, CLOY company records the following journal entries:

                                                                                       P                                P

Dr. Loan receivable                                                    1,500,000

        Cr. Cash                                                                                            1,500,000

(To recognize loan asset at gross amount)


Dr Impairment loss in profit or loss                              17,857

Cr Loss allowance in financial position                                          17,857

(To recognize 12-month expected credit losses)


4.). If there is no significant deterioration of the credit quality for the year ended 2021, then there will be no change to the recognition of the 12-month expected credit losses.

The journal entry for 12-month expected credit losses will be the same:

                                                                                             P                         P

Dr Impairment loss in profit or loss                                17,857

        Cr Loss allowance in financial position                                                17,857

(To recognize 12-month expected credit losses)


5.). If there is significant deterioration of credit risk, we will deduct the 12-month expected credit losses from the lifetime expected credit losses.

The lifetime expected credit losses = 72,096 – 17,857 = 54,239

The journal entry for the lifetime expected credit losses will be recorded as follows:

                                                                                             P                          P

Dr Impairment loss in profit or loss                                54,239

        Cr Loss allowance in financial position                                                54,239

(To recognize the lifetime expected credit losses)